The problem is 4x/5 + 3/2 = 2x the answer I got was x=1.25. I checked it with 1.25 in the equation and it works but all the problems I do like this seem to come out with decimal answers and I don't know if that is right. I feel like it's not. Help please!
Answer by rajagopalan(174)

solution:
4x/5 + 3/2 = 2x
multiply both sides by 10
8x + 15 = 20x
15=20x-8x
15=12x
15/12=x
1.25=x
your answer is perfectly ok
bravo